A stone of mass 100 g is suspended from one end of weightless string of length 1 m an d is allowed to swing in vertical plane .The speed of mass is 2 m/s when making angle of 60° with the vertical.Calculate the tension in the string when it is at 60°. And also calculate the velocity of body when it is at lowest position

Grade:12

1 Answers

Vikas TU
4 years ago
Dear student 
If you draw free body diagram you can see that T-mgcosa= mv​2/r
Solve for T
T=0.89
By using conservation of energy 
mgh1+mv1^2/2=mgh2+mv2^2/2
 Note: here work done by tension will not be considered since its direction of motion is                                                     perpendicular to force
mg/2=m(v2^2-v1^2)/2
v2=13.8^(1/2)